OpenStudy (anonymous):

How do I find if (1-5n^4)/((n^4)+(8n^3)) is convergent or not?

OpenStudy (anonymous):

As a sequence?

OpenStudy (anonymous):

Yes, as a sequence.

OpenStudy (anonymous):

You could consider it a function of x instead, and take the limit as x goes to infinity...

OpenStudy (anonymous):

Or how about this... \[\huge \frac{1-5n^4}{n^4+8n^3}\times\frac{\frac1{n^4}}{\frac1{n^4}}\]

OpenStudy (anonymous):

Basically multiplying it by 1 :)